# Entropy Rectifying Guidance for Diffusion and Flow Models

### v1

@@ Abstract
Guidance techniques are commonly used in diffusion and flow models to improve image quality and consistency for conditional generative tasks such as class-conditional and text-to-image generation. In particular, classifier-free guidance (CFG) — the most widely adopted guidance technique — contrasts conditional and unconditional predictions to improve the generated images. This results, however, in trade-offs across quality, diversity and consistency, improving some at the expense of others. While recent work has shown that it is possible to disentangle these factors to some extent, such methods come with an overhead of requiring an additional (weaker) model, or require more forward passes per sampling step. In this paper, we propose Entropy Rectifying Guidance (ERG), a simple and effective guidance mechanism based on inference-time changes in the attention mechanism of state-of-the-art diffusion transformer architectures, which allows for simultaneous improvements over image quality, diversity and prompt consistency. ERG is more general than CFG and similar guidance techniques, as it extends to unconditional sampling. ERG results in significant improvements in various generation tasks such as text-to-image, class-conditional and unconditional image generation. We also show that ERG can be seamlessly combined with other recent guidance methods such as CADS and APG, further boosting generation performance.
@@ 1 Introduction
Diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Dhariwal and Nichol, 2021) and flow models (Lipman et al., 2023; Ma et al., 2024; Esser et al., 2024) are state-of-the-art generative modeling tools for various modalities, ranging from images (Rombach et al., 2022; Podell et al., 2024; Chen et al., 2024; Esser et al., 2024), audio (Wang et al., 2023; Levy et al., 2023), and video (Jin et al., 2025; Polyak et al., 2024). These models generate data by starting with a simple prior and iteratively removing noise – a process called “denoising”. See Kingma and Gao (2023); Lipman et al. (2023) for details on diffusion and flow matching. These models can be conditioned on various inputs to control the generative process, e.g., in text-to-image models. Guidance techniques such as classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2021) are commonly used to improve sample quality and consistency with input conditioning. In the sampling process, these techniques combine a conditional signal with an unconditional one, and control the strength of conditioning through a scaling parameter. Although such guidance techniques are a crucial component in achieving state-of-the-art results, they are also known to affect the diversity of generated samples given a particular prompt (Sadat et al., 2024; Karras et al., 2024; Kynkäänniemi et al., 2024; Saharia et al., 2022). Moreover, too high guidance scales may lead to overly saturated images, affecting the quality of generated images (Saharia et al., 2022). An extensive analysis on the trade-offs of image generation quality, diversity, and consistency was presented by Astolfi et al. (2024). To mitigate quality-diversity-consistency trade-offs, more advanced guidance techniques have recently been proposed (Sadat et al., 2025; Karras et al., 2024). Most guidance approaches require spending part of the training cycles on unconditional generation for the guidance to work well (Sadat et al., 2025; Ho and Salimans, 2021), even if unconditional generation is not a goal in itself. They often rely on another model with weaker performance than the main model, thereby increasing memory requirements (Karras et al., 2024), or are not applicable to unconditional image generation (Sadat et al., 2025; Ho and Salimans, 2021).
@@
In our work, we build upon the work of Karras et al. (2024) and Hong (2024) and propose a simple and effective method to obtain both a strong and a weak predictive signal from a single model that leverages attention layers, where the model may be conditional or unconditional. In particular, our method uses the Hopfield energy formulation of attention (Ramsauer et al., 2021) and applies a temperature scaling to the softmax function of the attention layers in order to obtain the weak predictive signal. This scaling does not require any adaptations in model training, and may be applied to pre-trained denoising models, as well as their accompanying text encoders. Moreover, motivated by this energy interpretation of attention layers (Hong, 2024; Ramsauer et al., 2021), we also consider iterative re-application of attention layers, and rescaling the residual attention update. Our experiments show that manipulating the attention layers in this manner results in simultaneous improvements in sample quality, diversity, and consistency, contrary to the trade-offs observed by Astolfi et al. (2024) for classifier-free guidance.
@@
    Experimentally, we find that our guidance method significantly improves image quality and diversity while retaining the same prompt consistency as the standard classifier-free guidance ($+30$ points in density, $+4$ points in VQAScore) on COCO@512 using our 1.9 text-to-image image.
@@ 2 Related work and background
In this section, we briefly review background material on diffusion models, and related work on guidance techniques to improve sampling as well as the Hopfield energy formulation of attention.
@@ 2.1 Diffusion and flow matching models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) form a flexible class of generative models whose underlying principle is to map samples $\boldsymbol{\epsilon}$ from a trivial unit Gaussian prior $p_{0}=\mathcal{N}(0,{\bf I})$ to samples from a learned model $p_{1}$ of the data distribution. The forward process is defined as:
${\bf x}_{t}=\alpha_{t}{\bf x}_{1}+\sigma_{t}\boldsymbol{\epsilon},\quad\textrm{ with }\quad t\in[0,1],$ (1)
where $x_{1}\sim p_{1}$, and $\alpha_{t}$ is a decreasing function of “time” $t$ while $\sigma_{t}$ an increasing function of $t$.
@@ 2.2 Guidance mechanisms
Classifier guidance.  To enable high quality conditional generation, Dhariwal and Nichol (2021) proposed to guide the sampling process by leveraging gradients from pre-trained auxiliary classifier $p(c|{\bf x})$ in each reverse denoising step. They use the classifier to define the (scaled) joint score function as:
$\nabla_{{\bf x}_{t}}\log p({\bf x}_{t},c)=\nabla_{{\bf x}_{t}}\log p({\bf x}_{t})+w\nabla_{{\bf x}_{t}}\log p(c|{\bf x}_{t}),$ (2)
@@
where $p({\bf x}_{t})$ is an unconditional data model, and $w$ is a scalar parameter regulating the strength of the classifier guidance. While classifier guidance allows to improve input consistency and image quality (at the expense of diversity), it requires an auxiliary classification model that is robust to inputs ${\bf x}_{t}$ with varying amounts of noise.
@@
Classifier-free guidance.  To avoid the need for an auxiliary noise-robust classifier, Ho and Salimans (2021) proposed classifier-free guidance (CFG). In this case, during the training process, two generative models are learned, one conditional $p({\bf x}|c)$ and one unconditional $p({\bf x}|\emptyset)$. In practice, the unconditional model is trained by dropping conditioning information with a small probability. The score function used for sampling is extrapolated towards the conditional prediction and away from the unconditional prediction:
$\nabla_{\bf x}^{\textrm{CFG}}\log p({\bf x}|c)=w\nabla_{\bf x}\log p({\bf x}|c)+(1-w)\nabla_{\bf x}\log p({\bf x}|\emptyset).$ (3)
@@
While CFG improves image quality and input consistency with respect to classifier guidance, it tends to come at the cost of a reduction in diversity (Astolfi et al., 2024; Ho and Salimans, 2021; Sadat et al., 2024). Moreover, CFG often leads to generation artifacts, such as over-saturation as the guidance scale grows (Sadat et al., 2025).
@@
Advanced guidance techniques.  Several improved variants of classifier-free guidance have been proposed recently. Hong et al. (2023) presented Self-Attention Guidance (SAG), a guidance mechanism based on feeding a modified intermediate sample ${\bf x}_{t}$ when performing inference for unconditional prediction. The modification consists in blurring ${\bf x}_{t}$ in regions that are most attended to by the model’s self-attention. This method has been developed for the UNet architecture, hence applying it to more recent diffusion transformer architectures requires a hyperparameter search to understand which attention layers should be used for this method.
@@
Smoothed energy guidance (SEG) (Hong, 2024) contrasts conditional prediction with a “weaker” conditional prediction obtained by altering the attention’s softmax energy with a Gaussian kernel applied to queries. This method is developed for U-Net-style architectures and the softmax alteration applies to the self-attention layers in the middle block of the U-Net. Note that SEG does not rely on the difference between conditional and unconditional terms, and applies to both conditional and unconditional sampling. In the conditional case, SEG uses a linear combination of the conditional, unconditional, and energy-smoothed unconditional prediction, whereas in our approach we only use the conditional and smoothed conditional term. Thus, SEG requires an additional function evaluation with respect to ERG for conditional inference. For unconditional inference, both approaches require only two function evaluations. In a similar spirit, Ahn et al. (2024) propose a guidance method based on manipulating the attention mechanism, by replacing the attention matrix with an identity mapping inside the denoiser U-Net.
@@
Karras et al. (2024) explore using a smaller/weaker version of the same conditional model for classifier-free guidance, resulting in better diversity and image quality. Like ours, their approach can also be applied to unconditional sampling. However, their method requires access to an earlier checkpoint of the model or, for best results, training a separate model with lower capacity, as well as accessing two models when sampling, which may increase the memory footprint by a factor two.
@@
Rather than considering modifications of the unconditional model term, Sadat et al. (2024) proposed the “condition-annealed diffusion sampler” (CADS) to increase the diversity of generations while maintaining sample quality. This is achieved by adding Gaussian noise to the conditioning tokens during inference, using a piecewise-linear decreasing schedule on the noise amplitude. Sadat et al. (2025) propose “adaptive projected guidance” (APG), a variant of classifier-free guidance that resolves the over-saturation problem by emphasizing guidance orthogonal to the conditional prediction, rescaling the guidance term, and introducing a negative momentum term. The latter two approaches can be easily combined with our approach, and we consider such combinations in our experiments.
@@
Lastly, several works (Kynkäänniemi et al., 2024; Pavasovic et al., 2025; Wang et al., 2024) explore non-constant weight schedules for CFG. Such methods are complementary to our work, as our method only operates at the architecture level by rectifying the attention updates. Although they therefore could be combined with our approach, we defer this to future work.
@@ 2.3 Hopfield energy formulation of attention
The Hopfield network (Hopfield, 1982) is a dense associative memory model that aims to associate an input with its most similar pattern. More specifically, it constructs an energy function to model an energy landscape that contains basins of attraction around the desired patterns. Modern Hopfield energy networks (Ramsauer et al., 2021) introduce a new family of energy functions that improve the storage capacity of the model and make it compatible with continuous embeddings. Specifically, the following energy functional matches a continuous $d$-dimensional state (query) pattern $\boldsymbol{\xi}\in\mathbb{R}^{d}$ with $N$ stored (key) patterns ${\bf X}=({\bf x}_{1},...,{\bf x}_{N})\in\mathbb{R}^{d\times N}$ as:
$E(\boldsymbol{\xi};{\bf X})=\frac{1}{2}\boldsymbol{\xi}^{\top}\boldsymbol{\xi}-\text{LogSumExp}\left({\bf X}^{\top}\boldsymbol{\xi},\beta\right),$ (4)
where $\text{LogSumExp}({\bf x},\beta)=\beta^{-1}\log\left(\sum_{i=1}\exp(x_{i})\right)$, where $x_{i}$ are the elements of the vector ${\bf x}$, and $\beta$ a scalar hyperparameter defining the sharpness of the approximation of the maximum in the LogSumExp operation. Intuitively, the first term imposes a finite norm on the queries while the second term measures the alignment between the state patterns (queries) and stored patterns (keys). Using the Concave-Convex Procedure (CCCP) (Yuille and Rangarajan, 2003), an iterative update rule, which converges to the global minima of the energy, can be derived as:
$\boldsymbol{\xi}_{l+1}={\bf X}\;\text{SoftMax}(\beta{\bf X}^{\top}\boldsymbol{\xi}_{l}),$ (5)
where $\text{SoftMax}({\bf x})=\exp({\bf x}-\text{LogSumExp}({\bf x},1))$. Equivalently, each iterative update step can be seen as a gradient update in the negative direction of the energy:
$\nabla_{\boldsymbol{\xi}}E(\boldsymbol{\xi};{\bf X})=\boldsymbol{\xi}-{\bf X}\;\text{SoftMax}(\beta{\bf X}^{\top}\boldsymbol{\xi}).$ (6)
Taking a gradient descent step on this energy landscape with a step size $\gamma$ results in an update of the form $\boldsymbol{\xi}_{l+1}=\boldsymbol{\xi}_{l}-\gamma\left(\boldsymbol{\xi}_{l}-{\bf X}\text{softmax}(\beta{\bf X}^{\top}\boldsymbol{\xi}_{l})\right)$, and for $\gamma=1$ we recover the CCCP update.
@@
Ramsauer et al. (2021) show that the CCCP update is related to the standard attention operation as follows. Assuming there are $S$ state (query) patterns, and $N$ stored (key) patterns that can be mapped to keys, queries and values using linear transformations, he state pattern can be obtained through a concatenation: $\boldsymbol{\Xi}=[\boldsymbol{\xi}_{1},\dots,\boldsymbol{\xi}_{S}]\in\mathbb{R}^{d\times S}$. Then the attention map is given by $\Xi_{l+1}={\bf X}\text{softmax}({\bf X}^{\top}\Xi_{l})$. The keys, queries and values are obtained via linear projections as ${\bf K}={\bf X}W_{K}^{\top}\in\mathbb{R}^{N\times d},\ {\bf V}={\bf X}W_{V}^{\top}\in\mathbb{R}^{N\times d}$ and ${\bf Q}=\boldsymbol{\Xi}_{l+1}W_{Q}^{\top}\in\mathbb{R}^{S\times d}$, respectively. By setting $\beta=\frac{1}{\sqrt{d}}$ and substituting this into Equation 5, we obtain:
${\bf Q}_{l+1}$ $=\text{SoftMax}(\frac{{\bf Q}{\bf K}^{\top}}{\sqrt{d}}){\bf K}$ (7)
$=\text{SoftMax}\left(\frac{{\bf Q}{\bf K}^{\top}}{\sqrt{d}}\right){\bf V}\left(W_{K}W_{V}^{-1}\right)^{\top}.$ (8)
Hence, we observe the equation of attention up to a linear transformation, which provides an interpretation of the attention mechanism through the Hopfield energy lens. Such a formulation has been used in the Energy Based Cross-Attention method (EBCA) (Park et al., 2024) for adaptive context control in order to incorporate additional contexts into the generative process of a conditional diffusion model.

### v2

@@ Abstract
Guidance techniques are commonly used in diffusion and flow models to improve image quality and input consistency for conditional generative tasks such as class-conditional and text-to-image generation. In particular, classifier-free guidance (CFG) is the most widely adopted guidance technique. It results, however, in trade-offs across quality, diversity and consistency: improving some at the expense of others. While recent work has shown that it is possible to disentangle these factors to some extent, such methods come with an overhead of requiring an additional (weaker) model, or require more forward passes per sampling step. In this paper, we propose Entropy Rectifying Guidance (ERG), a simple and effective guidance method based on inference-time changes in the attention mechanism of state-of-the-art diffusion transformer architectures, which allows for simultaneous improvements over image quality, diversity and prompt consistency. ERG is more general than CFG and similar guidance techniques, as it extends to unconditional sampling. We show that ERG results in significant improvements in various tasks, including text-to-image, class-conditional and unconditional image generation. We also show that ERG can be seamlessly combined with other recent guidance methods such as CADS and APG, further improving generation results.
@@ 1 Introduction
Diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Dhariwal and Nichol, 2021) and flow models (Lipman et al., 2023; Ma et al., 2024; Esser et al., 2024) are state-of-the-art generative modeling tools for various modalities, ranging from images (Rombach et al., 2022; Podell et al., 2024; Chen et al., 2024; Esser et al., 2024), to audio (Wang et al., 2023; Levy et al., 2023), and video (Jin et al., 2025; Polyak et al., 2024). These models generate data by starting with a simple prior and iteratively removing noise – a process called “denoising”. See Kingma and Gao (2023); Lipman et al. (2023) for details on diffusion and flow matching. These models can be conditioned on various inputs to control the generative process, e.g., in text-to-image models. Guidance techniques such as classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2021) are commonly used to improve sample quality and consistency with input conditioning. In the sampling process, these techniques combine a conditional signal with an unconditional one, and control the strength of conditioning through a scaling parameter. Although such guidance techniques are a crucial component in achieving state-of-the-art results, they are also known to negatively affect the diversity of generated samples given a particular prompt (Sadat et al., 2024; Karras et al., 2024; Kynkäänniemi et al., 2024; Saharia et al., 2022). Moreover, too high guidance scales may lead to overly saturated images, affecting the quality of generated images (Saharia et al., 2022). An extensive analysis on the trade-offs of image generation quality, diversity, and consistency was presented by Astolfi et al. (2024). To mitigate these quality-diversity-consistency trade-offs, more advanced guidance techniques have recently been proposed, see e.g. Sadat et al. (2025); Karras et al. (2024). Most guidance techniques , however, require spending part of the training cycles on unconditional generation for the guidance to work, even if unconditional generation is not a goal in itself, and are not applicable to unconditional sampling (Sadat et al., 2025; Ho and Salimans, 2021). Others rely on a second model with weaker performance than the main model, thereby increasing memory requirements (Karras et al., 2024).
@@
In our work, we build upon the work of Karras et al. (2024) and Hong (2024), and propose Entropy Rectifying Guidance (ERG): a simple and effective method to obtain both a strong and a weak predictive signal from a single model that leverages attention layers, where the model may be conditional or unconditional. In particular, our method uses the Hopfield energy formulation of attention (Ramsauer et al., 2021; Hong, 2024) and applies a temperature scaling to the softmax function of the attention layers in order to obtain the weak predictive signal. This scaling does not require any adaptations in model training, and may be applied to pre-trained denoising models, as well as their accompanying text encoders. Moreover, motivated by this energy interpretation of attention layers, we also consider iterative re-application of attention layers, and rescaling the residual attention update. Our experiments show that manipulating the attention layers in this manner results in simultaneous improvements in sample quality, diversity, and consistency, contrary to the trade-offs observed by Astolfi et al. (2024) for classifier-free guidance.
@@
    Experimentally, we find that ERG significantly improves image quality and diversity while retaining the same prompt consistency as the standard classifier-free guidance (+30+30 points in density, +4+4 points in VQAScore) on COCO at 512 resolution using our 1.9B text-to-image model.
@@ 2 Related work and background
In this section, we briefly review background material on diffusion and flow models, and related work on guidance techniques to improve sampling as well as the Hopfield energy formulation of attention.
@@ 2.1 Diffusion and flow matching models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) form a flexible class of generative models whose underlying principle is to map samples $\boldsymbol{\epsilon}$ from a trivial unit Gaussian prior $p_{0}=\mathcal{N}(0,{\bf I})$ to samples from a learned model $p_{1}$ of the data distribution. The forward process is defined as: ${\bf x}_{t}=\alpha_{t}{\bf x}_{1}+\sigma_{t}\boldsymbol{\epsilon}$ with $t\in[0,1]$, where ${\bf x}_{1}\sim p_{1}$, and $\alpha_{t}$ is a decreasing function of “time” $t$ while $\sigma_{t}$ an increasing function of $t$.
@@ 2.2 Guidance mechanisms
Classifier guidance.  To enable high quality conditional generation, Dhariwal and Nichol (2021) proposed to guide the sampling process by leveraging gradients from pre-trained auxiliary classifier $p(c|{\bf x})$ in each denoising step. They use the classifier to define the (scaled) joint score function as $\nabla_{{\bf x}_{t}}\log p({\bf x}_{t},c)=\nabla_{{\bf x}_{t}}\log p({\bf x}_{t})+w\nabla_{{\bf x}_{t}}\log p(c|{\bf x}_{t}),$ where $p({\bf x}_{t})$ is an unconditional data model, and $w$ is a scalar parameter regulating the strength of the classifier guidance. While classifier guidance allows to improve input consistency and image quality (at the expense of diversity), it requires an auxiliary classification model that is robust to inputs ${\bf x}_{t}$ with varying amounts of noise.
@@
Classifier-free guidance.  To avoid the need for an auxiliary noise-robust classifier, Ho and Salimans (2021) proposed classifier-free guidance (CFG). In this case, during the training process, two generative models are learned, one conditional $p({\bf x}|c)$ and one unconditional $p({\bf x}|\emptyset)$. In practice, the unconditional model is trained by dropping conditioning information with a small probability. The score function used for sampling is extrapolated towards the conditional prediction and away from the unconditional prediction
@@
$\nabla_{\bf x}^{\textrm{CFG}}\log p({\bf x}|c)=w\nabla_{\bf x}\log p({\bf x}|c)+(1-w)\nabla_{\bf x}\log p({\bf x}|\emptyset).$ While CFG improves image quality and input consistency with respect to classifier guidance, it tends to come at the cost of a reduction in diversity (Astolfi et al., 2024; Ho and Salimans, 2021; Sadat et al., 2024). Moreover, CFG often leads to generation artifacts, such as over-saturation as the guidance scale $w$ grows (Sadat et al., 2025).
@@
Advanced guidance techniques.  Several improved variants of classifier-free guidance have been proposed recently. Hong et al. (2023) presented Self-Attention Guidance (SAG), a guidance mechanism based on feeding a modified intermediate sample ${\bf x}_{t}$ when performing inference for unconditional prediction. The modification consists in blurring ${\bf x}_{t}$ in regions that are most attended to by the model’s self-attention. This method has been developed for the U-Net architecture, hence applying it to more recent diffusion transformer architectures requires a hyperparameter search to understand which attention layers should be used for this method.
@@
Smoothed energy guidance (SEG) (Hong, 2024) contrasts conditional prediction with a “weaker” conditional prediction obtained by altering the attention’s softmax energy with a Gaussian kernel applied to queries. This method is developed for U-Net-style architectures and the softmax alteration applies to the self-attention layers in the middle block of the U-Net. In the conditional case, SEG uses a linear combination of the conditional, unconditional, and energy-smoothed unconditional prediction, whereas in our approach we only use the conditional and smoothed conditional term. Thus, SEG requires an additional function evaluation with respect to ERG for conditional inference. For unconditional inference, both approaches require only two function evaluations. In a similar spirit, Ahn et al. (2024) propose a guidance method based on manipulating the attention mechanism, by replacing the attention matrix with an identity mapping inside the denoiser U-Net.
@@
Karras et al. (2024) proposed AutoGuidance, a method that uses a smaller/weaker version of the same conditional model for classifier-free guidance, resulting in better diversity and image quality. Like ours, their approach can also be applied to unconditional sampling. However, their method requires access to an earlier checkpoint of the model or, for best results, training a separate model with lower capacity, as well as accessing two models when sampling, which may increase the memory footprint.
@@
Rather than considering modifications of the unconditional model term, Sadat et al. (2024) proposed the “condition-annealed diffusion sampler” (CADS) to increase the diversity of generations while maintaining sample quality. This is achieved by adding Gaussian noise to the conditioning tokens during inference, using a piecewise-linear decreasing schedule on the noise amplitude. Sadat et al. (2025) propose “adaptive projected guidance” (APG), a variant of CFG that resolves the over-saturation problem by emphasizing guidance orthogonal to the conditional prediction, rescaling the guidance term, and introducing a negative momentum term. The latter two approaches can be easily combined with our approach, and we consider such combinations in our experiments.
@@
Lastly, several works (Kynkäänniemi et al., 2024; Chung et al., 2024b; Pavasovic et al., 2025; Wang et al., 2024) explore non-constant weight schedules for CFG. Such methods are complementary to our work, as our method only operates at the architecture level by rectifying the attention updates. Although they therefore could be combined with our approach, we defer this to future work.
@@ 2.3 Hopfield energy formulation of attention
The Hopfield network (Hopfield, 1982) is a dense associative memory model that aims to associate an input with its most similar pattern. More specifically, it constructs an energy function to model an energy landscape that contains basins of attraction around the desired patterns. Modern Hopfield energy networks (Ramsauer et al., 2021) introduce a new family of energy functions that improve the storage capacity of the model and make it compatible with continuous embeddings. Specifically, the following energy functional matches a continuous $d$-dimensional state (query) pattern $\boldsymbol{\xi}\in\mathbb{R}^{d}$ with $N$ stored (key) patterns ${\bf X}=({\bf x}_{1},...,{\bf x}_{N})\in\mathbb{R}^{d\times N}$ as $E(\boldsymbol{\xi};{\bf X})=\frac{1}{2}\boldsymbol{\xi}^{\top}\boldsymbol{\xi}-\text{LogSumExp}\left({\bf X}^{\top}\boldsymbol{\xi},\beta\right),$ where $\text{LogSumExp}({\bf x},\beta)=\beta^{-1}\log\left(\sum_{i=1}^{d}\exp(x_{i})\right)$, where $x_{i}$ are the elements of the vector $\bf x$}, and $\beta$ a scalar hyperparameter defining the sharpness of the approximation of the maximum in the LogSumExp operation. Intuitively, the first term imposes a finite norm on the queries while the second term measures the alignment between the state patterns (queries) and stored patterns (keys). Using the Concave-Convex Procedure (CCCP) (Yuille and Rangarajan, 2003), an iterative update rule, which converges to the global minima of the energy, can be derived as $\boldsymbol{\xi}_{l+1}={\bf X}\;\text{SoftMax}(\beta{\bf X}^{\top}\boldsymbol{\xi}_{l}),$ where $\text{SoftMax}({\bf x})=\exp({\bf x}-\text{LogSumExp}({\bf x},1))$. Equivalently, each iterative update step can be seen as a gradient update in the negative direction of the energy $\nabla_{\boldsymbol{\xi}}E(\boldsymbol{\xi};{\bf X})=\boldsymbol{\xi}-{\bf X}\;\text{SoftMax}(\beta{\bf X}^{\top}\boldsymbol{\xi}).$ Taking a gradient descent step on this energy landscape with a step size $\gamma$ results in an update of the form $\boldsymbol{\xi}_{l+1}=\boldsymbol{\xi}_{l}-\gamma\left(\boldsymbol{\xi}_{l}-{\bf X}\text{SoftMax}(\beta{\bf X}^{\top}\boldsymbol{\xi}_{l})\right)$, and for $\gamma=1$ we recover the CCCP update.
@@
Ramsauer et al. (2021) show that the CCCP update is related to the standard attention operation as follows. Assuming there are $S$ state (query) patterns, and $N$ stored (key) patterns that can be mapped to keys, queries and values using linear transformations, the state pattern can be obtained through a concatenation: $\boldsymbol{\Xi}=[\boldsymbol{\xi}_{1},\dots,\boldsymbol{\xi}_{S}]\in\mathbb{R}^{d\times S}$. Then the attention map is given by $\Xi_{l+1}={\bf X}\text{SoftMax}({\bf X}^{\top}\Xi_{l})$. The keys, queries and values are obtained via linear projections as ${\bf K}={\bf X}W_{K}^{\top}\in\mathbb{R}^{N\times d},\ {\bf V}={\bf X}W_{V}^{\top}\in\mathbb{R}^{N\times d}$ and ${\bf Q}=\boldsymbol{\Xi}_{l+1}W_{Q}^{\top}\in\mathbb{R}^{S\times d}$, respectively. By setting $\beta=\frac{1}{\sqrt{d}}$ and substituting this into the CCCP update rule, we obtain: ${\bf Q}_{l+1}=\text{SoftMax}(\frac{{\bf Q}{\bf K}^{\top}}{\sqrt{d}}){\bf K}=\text{SoftMax}\left(\frac{{\bf Q}{\bf K}^{\top}}{\sqrt{d}}\right){\bf V}\left(W_{K}W_{V}^{-1}\right)^{\top}.$ Hence, we observe the equation of attention up to a linear transformation, which provides an interpretation of the attention mechanism through the Hopfield energy lens. Such a formulation has been used in the Energy Based Cross-Attention method (EBCA) (Park et al., 2024) for adaptive context control in order to incorporate additional contexts into the generative process of a conditional diffusion model.